# Calculation of the total 10-th order QED contribution to the lepton magnetic moments

### v1

@@ II RESULTS IN DETAIL
We extract $a_{e}$ from QED Feynman graphs with $N_{l}=2$, $N_{\gamma}=1$, where we denote by $N_{l}$ and $N_{\gamma}$ the number of external lepton and photon lines in the graph; each graph may contain electron, muon, tau-lepton lines. We also assume that all graphs are one-particle irreducible and have no odd lepton loops (Furry’s theorem). Undirected graphs are used everywhere in the calculations.
@@
Table 1: The contributions of all gauge-invariant classes containing lepton loops to $A_{1}^{(10)}$, calculated with different random number generators
@@
Table 2: The contributions of all gauge-invariant classes without lepton loops to $A_{1}^{(10)}$, calculated with different random number generators
@@ III REDUCTION TO FINITE INTEGRALS
The method of reduction of $A_{1}^{(2n)}$ to finite integrals is described in detail in S. Volkov (2024). It gives one Feynman parametric integral for each Feynman graph. The final value is simply the sum of these integrals. To make each integral finite, a subtraction procedure with linear operators applied to the Feynman amplitudes of UV-divergent subgraphs is utilized before arriving at the Feynman parameters. We focus here only on the aspects that are important for the current implementation.
@@
In old calculations we used $M^{2}=m^{2}$. Here we use $M^{2}=-m^{2}$; this slightly reduces inter-graph cancellations; see Section V and Volkov (2023).
@@ V TECHNICAL INFORMATION AND NUMERICAL CANCELLATIONS
In our approach, all divergences are numerically cancelled at the level of the integrands. This cancellation leads to round-off errors. We evaluate the integrand values using interval arithmetic to control the errors. The samples that result in too large intervals are evaluated with increased precisions.

### v2

# Calculation of the total 10-th order QED contribution to the electron magnetic moment

@@ II RESULTS IN DETAIL
We extract $a_{e}$ from QED Feynman graphs with $N_{l}=2$, $N_{\gamma}=1$, where we denote by $N_{l}$ and $N_{\gamma}$ the number of external lepton and photon lines in the graph; each graph may contain electron, muon, tau-lepton lines. We also assume that all graphs are one-particle irreducible and have no odd lepton loops (Furry’s theorem). Since the contribution of a QED Feynman graph does not depend on the arrow directions, we use undirected graphs everywhere in the calculations. By the number of Feynman graphs we always mean the number of undirected graphs.
@@
Table 1: The contributions of all gauge-invariant classes containing lepton loops to $A_{1}^{(10)}$, calculated with different random number generators; by $N_{\text{graph}}$ we denote the number of undirected Feynman graphs
@@
Table 2: The contributions of all gauge-invariant classes without lepton loops to $A_{1}^{(10)}$, calculated with different random number generators; by $N_{\text{graph}}$ we denote the number of undirected Feynman graphs
@@ III REDUCTION TO FINITE INTEGRALS
The method of reduction of $A_{1}^{(2n)}$ to finite integrals is described in detail in S. Volkov (2024). It gives one Feynman parametric integral for each Feynman graph. The final value is simply the sum of these integrals. To make each integral finite, a subtraction procedure with linear operators applied to the Feynman amplitudes of ultraviolet (UV) divergent subgraphs is utilized before arriving at the Feynman parameters. The use of direct subtraction of divergences under the integral sign is very important at the 5-loop level: the amount of computer resources required would be otherwise astronomically large. We focus here only on the aspects that are important for the current implementation.
@@
The Feynman graphs contributing to the anomalous magnetic moment contain ultraviolet and infrared (IR) divergences (and the mixed ones). All divergences cancel out in the final result since the on-shell renormalization is applied. However, any direct BPHZ-like implementation of the on-shell renormalization leaves individual graphs IR-divergent. To remove the IR divergences, the combinatorics of the subtraction must be modified (it is explained in detail in S. Volkov (2024)). The operators $U$ are used to remove UV divergences without the generation of additional IR divergences (and sometimes to remove IR divergences without generating additional UV divergences). All contributions of the terms containing $U$-operators are cancelled in the final result. In reality, the $U$-operators only redistribute the divergences between different graphs and regions in Feynman parametric space, this is why we have a free hand in the definition of the $U$-operators.
@@
The preservation of the Ward identity plays an important role in the equivalence of the subtraction method used with the on-shell renormalization: If $\Sigma$ and $\Gamma$ satisfy
$\Gamma_{\mu}(p,0)=-e\frac{\partial\Sigma(p)}{\partial p^{\mu}},$
where $e$ is the lepton charge, then
$(U\Gamma)_{\mu}(p,0)=-e\frac{\partial(U\Sigma)(p)}{\partial p^{\mu}}$
is also satisfied. In old calculations we used $M^{2}=m^{2}$. Here we use $M^{2}=-m^{2}$; this slightly reduces inter-graph cancellations; see Section V and Volkov (2023).
@@
The divergences associated with the photon-photon scattering subgraphs are removed by subtracting the value at zero momenta.
@@ V TECHNICAL INFORMATION AND NUMERICAL CANCELLATIONS
In our approach, all divergences are numerically cancelled at the level of the integrands. This cancellation leads to round-off errors. We evaluate the integrand values using interval arithmetic to control the errors. The samples that result in too large intervals are evaluated with increased precisions. The algorithm of interval acceptance and rejection was described in detail in S. Volkov (2019) (Section IV.A). It is constructed so that the round-off error is small relative to the Monte Carlo statistical error even in the worst case and takes into account that the round-off error can have a non-zero mean value and we cannot add them as statistical errors. In this calculation, it was modified so that it also provides correct values for the needed graph subsets.
@@
We evaluate all Feynman graphs directly; we never use known lower-order formulas for the vacuum polarization. These substitutions could significantly improve the precision for classes such as Set I(a), but the impact on the overall accuracy is negligible, and the individual precision achieved is enough for verification. For example, it is dominated by Set V that has no vacuum polarizations at all [^6] There are many graphs that have only one 1-loop vacuum polarization insertion. They provide a significant contribution (if we ignore Set V), but we do not believe that using explicit one-loop formulas would significantly improve the speed and precision: The dimensionality of the integrand would be reduced by 2 (from 13 to 11), but the integrand structure would become more complicated.. On the other hand, the use of explicit formulas makes the code more complicated and unreliable.